OpenStudy (anonymous):

As it turned out, Avery spent 5 hours per day driving and drove a total of 5,565 miles. What was his average speed for the hours he drove?

OpenStudy (anonymous):

5hours per day (5*7=35) 5565/35=159 159/5=31